A company tested 735 of the lightbulbs they produced and found them to have a mean life of 1,200 hours and a standard deviation of 50 hours. How many of these lightbulbs had a life between 1,170 hours and 1,230 hours?


1
Expert's answer
2021-01-13T17:01:18-0500

"P(1170\le X \le1230)=\\
P(\frac{1170-1200}{50}\le Z \le\frac{1230-1200}{50})=\\
F(\frac{1230-1200}{50})-F(\frac{1170-1200}{50})=\\
F(0.6)-F(-0.6)=\\
0.7257 -0.2743 =\\
0.4514"